OpenStudy (anonymous):

The measure of an interior angle of a regular polygon is 165.6°. How many sides make up the regular polygon?

OpenStudy (maheshmeghwal9):

Regular Polygon's sum of interior angles is \[180^o(n-2)=Sum.\] Where n= number of sides of the regular polygon. Therefore according to ur question this equation must be 100% true \[180^o(n-2)=165.6^o n\] Thus find 'n' & u will get ur answer.

OpenStudy (anonymous):

sides = 360/exterior angle